# Masked Autoencoder Self Pre-Training for Defect Detection in Microelectronics

### v1

@@ 5.3 Defect Detection
For inline automated optical inspection in microelectronics manufacturing, defect detection is a common practice. We perform defect detection given an input image using the $\Delta B_{max}$ values given by TTA as labels. While an LED with $\Delta B_{max}$ larger than $20\%$ is classified as faulty, we predict the scalar $\Delta B_{max}$ values directly to have a continuous measure for LED quality. That is, having the model predict $\Delta B_{max}$ not only indicates whether the given LED is functional, but also how far the LED is from failure. We compare ViT pre-trained with our MAE-based framework for 1.000 epochs against purely supervised ViT and ViT pre-trained with ImageNet. Also, we compare our models against large standard CNN backbones like ResNet [17], EfficientNet [36] or VGG [34] as well as against recent CNN architectures specifically designed for defect detection in microelectronics. In particular, we evaluate the model by Zippelius et al. [52] intended for SAM images of solder joints as well as the model by Zhang et al. [48] intended for X-ray images of solder joints. All models are trained for up to 200 epochs.
We find that ViT pre-trained with our self pre-training framework outperforms all other CNN and transformer-based models, achieving an MSE improvement of $5.30\%$ compared to the best-performing baseline (see Table 4). When it comes to other ViT models, our model outperforms purely supervised ViTs by $21.1\%$ and ViTs pre-trained on Imagenet by $10.2\%$. Notably, while the best results are achieved by the largest self pre-trained transformer model, ViT-Ti already outperforms considerably larger baselines by $3.20\%$ with only 5.7 million parameters. Thus, ViT self pre-trained with MAE not only substantially outperforms other ViT models and state-of-the-art CNN defect detection models, but is already able to do so using the smallest available architecture size.
Table 4: Defect Detection. All models are trained for 200 epochs on the SAM dataset. We report mean squared error (MSE) scores on the test set, where a lower score is better. Also, we report the percental differences to the best model.
Model Method Params MSE $\mathrm{\Delta}$ to Best
SAM-CNN [52] supervised 4.2 M 0.0336 10.5%
XRAY-CNN [48] supervised 8.7 M 0.0337 10.9%
ResNet50 [17] supervised 23.5 M 0.0339 11.5%
EfficientNet-B7 [36] supervised 63.8 M 0.0331 8.9%
VGG-11 [34] supervised 128.8 M 0.0320 5.3%
ViT-Ti [13] supervised 5.7 M 0.0380 25.0%
ViT-S [13] supervised 22.1 M 0.0368 21.1%
ViT-B [13] supervised 86.7 M 0.0380 25.0%
@@
We perform interpretability analysis with class activation maps using GradCAM [32]. In GradCAM, the relevance of image regions for the output is visualized by computing gradients of the output with respect to a given layer. We choose the first norm layer of the last transformer block for the gradient computation. Results for all ViT-B variants are shown in Figure 5. We find that:
@@
    Supervised ViT-B has scattered attention for both early and late TSC (see Figure 5), indicating that the transformer encoder does not learn meaningful representations in purely supervised training.
@@
    ViT-B pre-trained on ImageNet strongly focuses on the bright bounding regions between solder pads, indicating that the model attends to the LED type, even when clear defects such as cracks in the solder material are visible (see Figure 5 (b), row 3 and 4). We interpret this as the model learning a shortcut rather than the actual dependency between defects like cracks in the solder material and $\Delta B_{max}$.
@@
    ViT-B self pre-trained on SAM data solely attends to actual defects such as cracks for later TSC with stronger degradation (Figure 5 (b)) and attends to voids and the boundary regions between solder pads if no damages are visible (Figure 5 (a)). This indicates that the model prioritizes actual defects and considers the general error-proneness of different LED types when no defects are visible.
@@
(b) 1.000 - 1.500 TSC
Figure 5: Attention Visualization. Compared to supervised ViT-B and ViT-B pre-trained with MAE on ImageNet, our model self pre-trained with MAE focuses on the LED type and on voids in the solder material for earlier TSC (a), and on defects such as cracks for later TSC (b), indicating that MAE self pre-training yields superior and defect-specific feature representations.
@@ 6 Discussion and Conclusion
The demonstrated effectiveness of our approach has broad implications for quality control in Industry 4.0 settings, where maximizing production yield while minimizing the cost and time required for inspection remains a central challenge. By enabling more accurate detection of developing solder joint defects in early stages, our approach can reduce warranty costs, improve product reliability, and increase customer satisfaction. As manufacturing processes continue to demand higher precision and reliability while dealing with ever-increasing miniaturization, approaches like ours that can leverage limited domain data effectively will become increasingly valuable. The combination of high accuracy, interpretability, and resource efficiency positions our MAE self pre-training framework as a promising path forward for next-generation quality inspection systems in microelectronics, with particular potential for integration into capable smart sensors that enable real-time defect detection directly on the production line.

### v2

@@ 5.3 Defect Detection
For inline automated optical inspection in microelectronics manufacturing, defect detection is a common practice. We perform defect detection given an input image using the $\Delta B_{max}$ values given by TTA as labels. While an LED with $\Delta B_{max}$ larger than $20\%$ is classified as defect, we predict the scalar $\Delta B_{max}$ values directly to have a continuous measure for LED quality. That is, having the model predict $\Delta B_{max}$ not only indicates whether the given LED is functional, but also how far the LED is from failure. Note that defect detection differs from anomaly detection where a pixel-wise labelling is performed (see e.g. [9]).
We compare ViT pre-trained with our MAE-based framework for 1,000 epochs against purely supervised ViT and ViT pre-trained with ImageNet. Also, we compare our models against large standard CNN backbones like ResNet [18] or EfficientNet [31] without additional pre-training, as well as against recent CNN architectures specifically designed for defect detection in microelectronics. In particular, we evaluate the model by Zippelius et al. [42] intended for SAM images of solder joints as well as the model by Zhang et al. [38] intended for X-ray images of solder joints. CNN models are trained in supervised fashion on defect detection without additional pre-training. All models are trained for up to 200 epochs.
We find that ViT pre-trained with our self pre-training framework outperforms all CNN and transformer-based models, achieving an MSE improvement of $8.9\%$ compared to the best-performing baseline (see Table 4). When it comes to other ViT models, our model outperforms purely supervised ViTs by $21.1\%$ and ViTs pre-trained on Imagenet by $10.2\%$. Notably, while the best results are achieved by the largest self pre-trained transformer model, ViT-Ti already outperforms considerably larger baselines by $6.9\%$ with only 5.7 million parameters. Thus, ViT self pre-trained with our approach not only substantially outperforms other ViT models and state-of-the-art CNN defect detection models, but is already able to do so using the smallest available architecture size.
Table 4: Defect Detection. All models are trained for 200 epochs on the SAM dataset. We report absolute MSE and relative difference compared to the best-performing model ($\mathrm{\Delta}$ to Best) to provide an intuitive sense of performance gains.
Model Pre-Training Params MSE $\mathrm{\Delta}$ to Best
SAM-CNN [42] - 4.2 M 0.0336 10.5%
XRAY-CNN [38] - 8.7 M 0.0337 10.9%
ResNet50 [18] - 23.5 M 0.0339 11.5%
EfficientNet-B7 [31] - 63.8 M 0.0331 8.9%
ViT-Ti [14] - 5.7 M 0.0380 25.0%
ViT-S [14] - 22.1 M 0.0368 21.1%
ViT-B [14] - 86.7 M 0.0380 25.0%
@@
5.3.1 Interpretability Analysis
We analyze class activation maps using GradCAM [30]. In GradCAM, the relevance of image regions for the output is visualized by computing gradients of the output with respect to a given layer. We choose the first norm layer of the last transformer block for the gradient computation. Results for all ViT-B variants are shown in Figure 5. We find that:
@@ 5.3.1 Interpretability Analysis
    Supervised ViT-B has scattered attention for both early and late TSC (see Figure 5). This suggests that the model fails to form semantically meaningful representations of defects under purely supervised training.
@@
    ViT-B pre-trained on ImageNet shows sharp but misleading focus — attending to the boundary regions around solder pads even when defects are present. These regions are not causally linked to degradation. The model thus appears to exploit spurious correlations involving the general lifetime of certain LED types, compromising its generalizability to unseen defect modes.
@@
    ViT-B self pre-trained on SAM data, in contrast, displays the behavior we expect from a robust and generalizable model:
    • For severe degradation cases (Figure 5 (b)), attention is focused on real defects such as cracks.
    • For defect-free samples (Figure 5 (a)), attention shifts to error-prone areas such as void zones, or the bounding areas which indicate the type of the given LED, which is related to its general life expectancy.
@@
(b) 1,000 - 1,500 TSC
Figure 5: Attention Visualization. GradCAM heatmaps for supervised ViT, ViT pre-trained on ImageNet, and our ViT self pre-trained on SAM data. Baseline models have scattered attention or learn shortcuts between LED type and defects, thus neglecting the true causal relationship between damages in the solder and LED defects. Our self-pre-trained model, in contrast, consistently focuses on general structural information when no defects are present (a), and pinpoints defect-relevant regions such as cracks whenever they are visible (b).
@@ 6 Discussion and Conclusion
While our MAE-based self pre-training framework demonstrates strong performance in defect detection, several directions offer potential for further enhancement. Future work could explore multi-modal pre-training that integrates additional signals such as thermal profiles or X-ray imaging to enrich learned representations. Combining images of multiple timesteps during pre-training may enable earlier and more accurate failure forecasting. Moreover, adaptive or defect-aware masking strategies could guide the model toward more semantically relevant features. Extending the framework to other manufacturing domains or unseen microelectronic devices would also test its robustness and generalizability. Finally, integrating our models into edge-computing environments for real-time, in-situ quality control remains a critical step toward practical deployment.